# UniTacHand: Unified Spatio-Tactile Representation for Human to Robotic Hand Skill Transfer

### v2

@@ 1 Introduction
Dexterous manipulation for contact-rich tasks remains a grand challenge in robotics [34, 40, 24, 18], largely because the systems lack the rich, multi-modal tactile perception that humans use. While vision-based policies [5, 54, 48] are dominant, they struggle with fundamental limitations like occlusion and ambiguity in complex interactions [21, 49]. Tactile sensing provides a distinct, complementary information stream that addresses these shortcomings [6]. However, despite the advantage of diverse, high-resolution tactile sensors, effectively processing and unifying heterogeneous tactile data remains an open problem [10].
To endow robots with autonomous manipulation capabilities, data-driven methods have become increasingly prominent. Reinforcement learning (RL) for robotics, while showing promise for some tasks, often suffers from high sample complexity [16] and the well-known sim-to-real gap – especially pronounced in tasks with complex contact dynamics [1, 3, 15]. Imitation learning (IL) and learning from demonstration (LfD) methods have emerged as a promising alternative, directly mimicking expert data to acquire manipulation skills. However, applying IL to tactile-based dexterous manipulation is fundamentally limited by the general scarcity and low quality of real robot teleoperation data [52, 50]. LfD methods [14, 4, 28], especially those that utilize human data collected by wearable devices, are hindered by the embodiment gap between human and robotic hands [38].
Current efforts to bridge the embodiment gap have largely focused on kinematic retargeting [46, 14], leaving the tactile morphological gap between different robotic hands largely unaddressed. A common workaround is to equip human hands with the same robotic tactile sensors used on the robot for data collection [37, 32]. However, such approaches rely on hardware uniformity, which can be difficult to achieve for many robotic hands and tends to be less cost-effective or scalable for data collection. In this research, we aim to use human data collected by wearable haptic gloves to learn tactile-based robotic manipulation policies. Our focus thus shifts to the following question: how can we design a method to transfer tactile manipulation knowledge from haptic gloves to dexterous robotic hands?
@@
To bridge such cross-embodiment gaps, leveraging a shared canonical representation is essential. In the realm of kinematics and vision, the MANO hand model [31] has been established as a cornerstone, providing a parametric space that unifies diverse hand shapes and poses [12]. MANO model has been widely adopted as a general representation for learning policies on dexterous hands [36, 41, 46, 29, 42]. We argue that this geometric consistency can be extended beyond kinematics to unify tactile perception.
@@ 2 Related Work
Our work is built upon the foundation of parametric 3D hand models. Loper [27] first introduced SMPL, a learned parametric model for human bodies. Building on this, Romero [31] proposed MANO, a powerful and expressive model of hand shape and articulation derived from thousands of high-resolution 3D scans. MANO has become a standard in the computer vision community, which is widely employed for visual hand pose and shape estimation from images or videos [2, 51, 47]. Besides, as a kind of motion representation, MANO is also widely used in motion generative models and even vision-language-action (VLA) models [56, 28, 22, 8, 41]. However, its potential as a canonical coordinate frame for tactile data has been largely unexplored. Prior work may project contact onto a mesh [35]. Our work is the first to leverage the intrinsic, morphologically consistent MANO UV map as a unified spatio-tactile representation space specifically for cross-domain human-to-robot policy transfer.
@@
The diversity of tactile sensors (e.g., visual-tactile sensors [49, 19]) poses a significant domain adaptation challenge. Early efforts to learn a shared latent space [9, 23], typically via auto-encoders, discarded the precise spatial geometry essential for manipulation policies. Recent works seek unified representations, but diverge in their target alignment space. UniTouch [44] aligns to the semantic space of CLIP [30] for zero-shot semantic tasks. Some other works including T3 [53] use physics-based priors (e.g., forces [39]) or learn data-hungry implicit embeddings, which require millions of samples. On the other hand, paired video-tactility data have been leveraged to align force information with visual contact, extracting knowledge and skills from human demonstrations [33, 42]. UniTacHand presents a more data-efficient and geometrically grounded alternative. We utilize such spatial context as useful signals, and align it to an explicit morphological space. By leveraging the MANO UV map as a powerful inductive bias, our model learns a aligned embedding from a fraction of the data (a 10-minute collected paired set).
@@
Learning from demonstration (LfD) has been well-established for visual or kinematic trajectories [20, 1, 55, 28], yet transferring contact-rich policies via tactile data remains an open problem. The challenge stems from the embodiment gap – the profound morphological and sensory mismatch between a human hand and a robotic end effector. For instance, multiple human demonstration datasets are widely used to learn behaviors from human manipulation videos or trajectories [7, 17, 25, 11, 26], as such datasets provide strong priors of performing manipulation tasks [56, 28, 43].
In the field of tactile sensing, prior research has often tailored its approach according to the morphological complexity of the target system. For instance, Feel the Force (FTF) [32] achieved compelling zero-shot transfer by unifying representations of contact forces; however, this was demonstrated only on a parallel-jaw gripper, a simple structure that circumvents the challenges of high-DoF, multi-finger coordination. When it comes to more complex, anthropomorphic hands, MimicTouch [45] relies on a sophisticated three-stage pipeline that includes online residual reinforcement learning. This dependence on an online RL patch to correct the policy on the physical robot reveals that the embodiment gap has not been fundamentally addressed at the representation level. In contrast, our work focuses on resolving the embodiment gap by aligning different domains into a canonical space. This approach enables genuine zero-shot transfer for dexterous hands and enriches the learned knowledge by integrating human demonstrations with one-shot data from real robots.
@@ 3.1 Unified Representation via MANO UV Maps
The MANO model [31] provides a low-dimensional, differentiable parameterization for the human hand, defined by shape parameters $\beta$ and pose parameters $\theta$. Critically, it offers a consistent mesh topology ($N_{v}=778$ vertices) and a corresponding 2D UV map, which we utilize as our canonical surface space for decoders.
@@
To focus the learning process only on active sensing regions, we then create binary masks $M_{H}$ and $M_{R}$. These masks are 11 for texels that correspond to a sensor region and 0 otherwise. The final unified representations, which are used in all subsequent steps, are defined as:
@@ 4.1 Experimental Setup
During representation learning, we train the alignment encoders using a small, 10-minute paired dataset collected from both human and robotic dexterous hands simultaneously figure˜6. During data collection, we use retargeting algorithms including DexPilot [12] to map MANO keypoints captured from human hand to the Inspire dexterous hand by teleoperation, ensuring that both human and the robot hand are performing the same action in a similar behavioral manner. We collect such paired data on 50 kinds of diverse paired objects , performing the same manipulation process including grasping simultaneously with various hand poses. We collect 688 trajectories containing 16k frames in total, at a frequency of 40 Hz.
@@ A.1 Data Unification and Preprocessing
We utilize the standard MANO model [31] featuring $N_{v}=778$ vertices and $N_{f}=1538$ faces. All tactile projections are rasterized into a normalized 2D UV map with a resolution of $W\times H=1024\times 1024$.
@@ C.2 Training Policies for Tasks
We list the implementation details of the four policies (networks and architectures) in table˜5. For our one-shot manipulation task, we utilize a pre-trained ResNet-18 [13] as the visual backbone, and concatenate the visual embedding with the tactile embedding, which is then fed into the policy network.

### v3

@@ 1 Introduction
Dexterous manipulation for contact-rich tasks remains a grand challenge in robotics [suh2025dexterous, wu2021learning, liang2025dexhanddiff, lakshmipathy2024contactmpc], largely because the systems lack the rich, multi-modal tactile perception that humans use. While vision-based policies [diffusion-policy, dexgraspvla, demograsp] are dominant, they struggle with fundamental limitations like occlusion and ambiguity in complex interactions [levine2018learning, yuan2017gelsight]. Tactile sensing provides a distinct, complementary information stream that addresses these shortcomings [dahiya2010tactile]. However, despite the advantage of diverse, high-resolution tactile sensors, effectively processing and unifying heterogeneous tactile data remains an open problem [gao2021survey].
To endow robots with autonomous manipulation capabilities, data-driven methods have become increasingly prominent. Reinforcement learning (RL) for robotics, while showing promise for some tasks, often suffers from high sample complexity [kober2013reinforcement] and the well-known sim-to-real gap – especially pronounced in tasks with complex contact dynamics [argall2009survey, chebotar2019closing, resdex]. Imitation learning (IL) and learning from demonstration (LfD) methods have emerged as a promising alternative, directly mimicking expert data to acquire manipulation skills. However, applying IL to tactile-based dexterous manipulation is fundamentally limited by the general scarcity and low quality of real robot teleoperation data [aloha, doglove]. LfD methods [hoque2025egodex, chen2022dextransfer, beingh0], especially those that utilize human data collected by wearable devices, are hindered by the embodiment gap between human and robotic hands [wang2023cross].
Current efforts to bridge the embodiment gap have largely focused on kinematic retargeting [crossdex, hoque2025egodex], leaving the tactile morphological gap between different robotic hands largely unaddressed. A common workaround is to equip human hands with the same robotic tactile sensors used on the robot for data collection [wang2023dexumi, FTF]. However, such approaches rely on hardware uniformity, which can be difficult to achieve for many robotic hands and tends to be less cost-effective or scalable for data collection. In this research, we aim to use human data collected by wearable haptic gloves to learn tactile-based robotic manipulation policies. Our focus thus shifts to the following question: how can we design a method to transfer tactile manipulation knowledge from haptic gloves to dexterous robotic hands?
@@
To bridge such cross-embodiment gaps, leveraging a shared canonical representation is essential. In the realm of kinematics and vision, the MANO hand model [MANO] has been established as a cornerstone, providing a parametric space that unifies diverse hand shapes and poses [handa2020dexpilot]. MANO model has been widely adopted as a general representation for learning policies on dexterous hands [wang2023memahand, xie2024ms, crossdex, mandi2025dexmachina, xu2025dexcanvas]. We argue that this geometric consistency can be extended beyond kinematics to unify tactile perception.
@@ 2 Related Work
Our work is built upon the foundation of parametric 3D hand models. loper2015smpl first introduced SMPL, a learned parametric model for human bodies. Building on this, MANO proposed MANO, a powerful and expressive model of hand shape and articulation derived from thousands of high-resolution 3D scans. MANO has become a standard in the computer vision community, which is widely employed for visual hand pose and shape estimation from images or videos [boukhayma20193d, zhang2021handoccnet, being0]. Besides, as a kind of motion representation, MANO is also widely used in motion generative models and even vision-language-action (VLA) models [zhou2025megohand, beingh0, li2025scalable, fu2025handrawer, xie2024ms]. However, its potential as a canonical coordinate frame for tactile data has been largely unexplored. Prior work may project contact onto a mesh [sundaralingam2023contactmap]. Our work is the first to leverage the intrinsic, morphologically consistent MANO UV map as a unified spatio-tactile representation space specifically for cross-domain human-to-robot policy transfer.
@@
The diversity of tactile sensors (e.g., visual-tactile sensors [yuan2017gelsight, DIGIT]) poses a significant domain adaptation challenge. Early efforts to learn a shared latent space [li2020connecting], typically via auto-encoders, discarded the precise spatial geometry essential for manipulation policies. Recent works seek unified representations, but diverge in their target alignment space. UniTouch [yang2024unitouch] aligns to the semantic space of CLIP [CLIP] for zero-shot semantic tasks. Some other works including T3 [zhao2024transferable] use physics-based priors (e.g., forces [wu2024canonical]) or learn data-hungry implicit embeddings, which require millions of samples. On the other hand, paired video-tactility data have been leveraged to align force information with visual contact, extracting knowledge and skills from human demonstrations [song2025opentouch, xu2025dexcanvas]. UniTacHand presents a more data-efficient and geometrically grounded alternative. We utilize such spatial context as useful signals, and align it to an explicit morphological space. By leveraging the MANO UV map as a powerful inductive bias, our model learns a aligned embedding from a fraction of the data (a 10-minute collected paired set).
@@
Learning from demonstration (LfD) has been well-established for visual or kinematic trajectories [laskin2020curl, argall2009survey, bidexhd, beingh0], yet transferring contact-rich policies via tactile data remains an open problem. The challenge stems from the embodiment gap – the profound morphological and sensory mismatch between a human hand and a robotic end effector. For instance, multiple human demonstration datasets are widely used to learn behaviors from human manipulation videos or trajectories [arctic, h2o, taco, grauman2022ego4d, liu2022hoi4d], as such datasets provide strong priors of performing manipulation tasks [zhou2025megohand, beingh0, yang2025egovla].
In the field of tactile sensing, prior research has often tailored its approach according to the morphological complexity of the target system. For instance, Feel the Force (FTF) [FTF] achieved compelling zero-shot transfer by unifying representations of contact forces; however, this was demonstrated only on a parallel-jaw gripper, a simple structure that circumvents the challenges of high-DoF, multi-finger coordination. When it comes to more complex, anthropomorphic hands, MimicTouch [yu2024mimictouch] relies on a sophisticated three-stage pipeline that includes online residual reinforcement learning. This dependence on an online RL patch to correct the policy on the physical robot reveals that the embodiment gap has not been fundamentally addressed at the representation level. In contrast, our work focuses on resolving the embodiment gap by aligning different domains into a canonical space. This approach enables genuine zero-shot transfer for dexterous hands and enriches the learned knowledge by integrating human demonstrations with one-shot data from real robots.
@@ 3.1 Unified Representation via MANO UV Maps
The MANO model [MANO] provides a low-dimensional, differentiable parameterization for the human hand, defined by shape parameters $\beta$ and pose parameters $\theta$. Critically, it offers a consistent mesh topology ($N_{v}=778$ vertices) and a corresponding 2D UV map, which we utilize as our canonical surface space for decoders.
@@
To focus the learning process only on active sensing regions, we then create binary masks $M_{H}$ and $M_{R}$. These masks are $1$ for texels that correspond to a sensor region and 0 otherwise. The final unified representations, which are used in all subsequent steps, are defined as:
@@ 4.1 Experimental Setup
During representation learning, we train the alignment encoders using a small, 10-minute paired dataset collected from both human and robotic dexterous hands simultaneously figure˜6. During data collection, we use retargeting algorithms including DexPilot [handa2020dexpilot] to map MANO keypoints captured from human hand to the Inspire dexterous hand by teleoperation, ensuring that both human and the robot hand are performing the same action in a similar behavioral manner. We collect such paired data on 50 kinds of diverse paired objects , performing the same manipulation process including grasping simultaneously with various hand poses. We collect 688 trajectories containing 16k frames in total, at a frequency of 40 Hz.
@@ A.1 Data Unification and Preprocessing
We utilize the standard MANO model [MANO] featuring $N_{v}=778$ vertices and $N_{f}=1538$ faces. All tactile projections are rasterized into a normalized 2D UV map with a resolution of $W\times H=1024\times 1024$.
@@ C.2 Training Policies for Tasks
We list the implementation details of the four policies (networks and architectures) in table˜5. For our one-shot manipulation task, we utilize a pre-trained ResNet-18 [resnet] as the visual backbone, and concatenate the visual embedding with the tactile embedding, which is then fed into the policy network.